OpenStudy (anonymous):

Given that lg x = p and lg y = q, express the following in terms of p and q. lg(y^x)

OpenStudy (anonymous):

y = E^q , x = E^q

OpenStudy (anonymous):

x = E^p rather

OpenStudy (anonymous):

Log((E^q)^(E^p)) use properties of logs E^p * Log[E^q] q*E^p